OpenStudy (anonymous):

19/sqr of 150 plus sqr of 3

OpenStudy (anonymous):

Is the +sqrt(3) in denominator?

jhonyy9 (jhonyy9):

\[19/(\sqrt150 +\sqrt3)=19/(5\sqrt6+\sqrt3)\]

OpenStudy (anonymous):

You might have to rationalize this, \[19/(\sqrt{150}+\sqrt{3})\times(\sqrt{150}-\sqrt{3})/(\sqrt{150}-\sqrt{3})\] \[(19\sqrt{150}-19\sqrt{3})/(150+3)\] \[(95\sqrt{6}-19\sqrt{3})/1\]